(4) Determine and draw the Thevenin equivalent circuit with respect to terminals a and b. 20ix i 452 42
Answer #1

Ini 3 a Th KCL at node u 4 m[ * ++++] - Mi[ 142+1) - 44,- V2 = 9 - KCL at node V. » V2 V1 + 3 x = 0 4 Since, V = -Us V2 -v1 +th=– 13 + 20 x 5 37 V I th= 37 Now Determine Rith: short circuit arb Find short erruit current Isc, and Uth Rth Isc & Isc 3vasubstitute above valun in 135-36 +3(-36) + Isc = 0 Isc = + 83.25 A Isc= 83.25 A Rith = Th = 3 - 83.25 = 0. = 0.444 n. Isc Rih